# Physics-Guided Multi-Fidelity DeepONet for Data-Efficient Flow Field Prediction

### v1

@@ 5.2 Performance Analysis with Limited High-Fidelity Data
The results shown in Figure 11 reveal several important insights about the MF-DeepONet’s performance with limited HF data. First, examining the MSE trends (Figure 11(a)), we observe a consistent pattern of improved prediction accuracy as the number of HF samples increases across all configurations. The 64×64 resolution configurations consistently outperform their 32×32 counterparts, with the 64×64 with 300 samples combination achieving the best performance across all HF dataset sizes. A particularly noteworthy observation is that increasing the LF dataset size from 200 to 300 samples yields more improvements in accuracy for the 64×64 resolution compared to the 32×32 resolution. For instance, with 60 HF samples, the 64×64 resolution shows a reduction in MSE from 13.476 (200 LF samples) to 11.364 (300 LF samples)—a 15.7% improvement—while the 32×32 resolution exhibits a smaller improvement from 16.031 to 13.760, representing a 14.2% reduction in error.
@@
Figure 11: MF-DeepONet performance with limited high-fidelity data: different resolutions of LF data with different LF data sizes are explored. The label “LF 32×32 (200)” represents a case where LF data is from a 32×32 resolution with a dataset size of 200—all other labels follow the same notation, indicating the corresponding LF resolution and dataset size.
The training time analysis (Figure 11(b)) reveals an expectable trade-off. Configurations with 300 LF samples consistently require longer training times compared to their 200 LF sample counterparts due to the longer training time required in LF deepONet. However, this increased computational cost appears justified by the improved accuracy, particularly for the 64×64 resolution cases.
6 Time-Derivative Guided Sampling for Enhancing Multi-Fidelity DeepONet
6.1 Effects of Time-Derivative Guided Sampling
Based on our analysis in Section 5.2, we now investigate two optimal configurations of our fine-tuned MF-DeepONet framework (highlighted in Figure 11(a) with yellow circles and black dashed edges) that achieve an exceptional balance between prediction accuracy and high-fidelity data efficiency:
@@ 6.1 Effects of Time-Derivative Guided Sampling
    Case 1: MF-DeepONet with 32×32 LF resolution (300 samples) trained using 80 HF samples
@@
    Case 2: MF-DeepONet with 64×64 LF resolution (300 samples) trained using 60 HF samples
@@
For these two configurations, we analyze the effectiveness of our time-derivative guided sampling strategy in practice. This physics-guided approach, formalized in Algorithm 1 in Section 2.2.2, strategically identifies regions with pronounced temporal dynamics from the pre-trained LF DeepONet, enabling more efficient HF data collection in physically meaningful areas.
In our experiments, we set the sampling ratio $r=0.1$, meaning that 10% of the HF spatial points are selected based on the temporal derivative information from LF DeepONet, while the remaining 90% are sampled uniformly at random to maintain spatial coverage. When compared to their respective baselines only with uniform sampling (without time-derivative guided sampling), both configurations show substantial improvements in prediction accuracy, as detailed in Table 5. The 32×32 LF resolution case shows a 10.02% reduction in MSE (from 11.759 to 10.581), while the 64×64 LF resolution case achieves a significant MSE reduction of 20.73% (from 11.364 to 9.008). These accuracy gains come with moderate increases in computational cost—7.68% and 10.74% increases in training time for Case 1 and Case 2, respectively. This additional computational overhead, primarily stemming from the automatic differentiation calculations required for temporal derivatives, is well justified given the significant improvement in accuracy. The greater improvement in the 64×64 configuration indicates that higher-resolution LF data produces a more accurate LF DeepONet model, which in turn provides more reliable temporal derivative information. This enhanced precision in identifying dynamically important regions results in more effective HF data sampling, improving the utilization of limited high-fidelity resources.
@@
Table 5: Performance comparison between uniform (Before) and time-derivative guided sampling strategies (After).
Metric Case Before After Change
MSE Case 1: LF 32×32 w/ 80 HF 11.759 10.581 -10.02%
Case 2: LF 64×64 w/ 60 HF 11.364 9.008 -20.73%
Training hour Case 1: LF 32×32 w/ 80 HF 0.469 0.505 +7.68%
Case 2: LF 64×64 w/ 60 HF 0.475 0.526 +10.74%
@@
The above quantitative improvements are further supported by the qualitative analysis shown in Figure 12, which compares vorticity profiles at different time steps. At $t=0.25s$ (Figure 12(a)), all models show similar prediction capabilities, closely following the ground truth profile. This comparable performance in the early stage is expected, as the flow field remains relatively simple and closer to the initial conditions. However, at $t=0.5s$ (Figure 12(b)), clear differences emerge in the models’ ability to capture complex flow dynamics. This later time step exhibits significantly more complicated flow development with pronounced temporal variations—precisely the conditions where our time-derivative sampling strategy should excel. Indeed, particularly in the middle region ($y=0.3m$ to $0.7m$), models trained with time-derivative sampling (presented as dashed lines) more accurately reproduce the local fluctuations and subtle variations in the vorticity profile. This superior performance demonstrates that our approach automatically identifies and focuses computational resources on regions with significant temporal dynamics, enhancing prediction accuracy exactly where conventional methods struggle. In contrast, models without sampling tend to oversimplify these local features, capturing only the overall trend while missing important flow structure details. These results confirm that our physics-guided sampling strategy effectively targets time-sensitive regions of the flow field ($t=0.5s$ than $t=0.25s$), showing particular advantage in predicting complex time-dependent dynamics that develop as the simulation progresses.
@@
Figure 12: Comparison of vorticity profiles from ground truth (GT) and MF-DeepONet predictions with and without time-derivative guided sampling. At early stages ($t=0.25s$), all models show comparable performance, but at later times ($t=0.5s$), models with time-derivative sampling demonstrate superior accuracy in capturing complex flow dynamics. Case 1: 32×32 LF (300 samples) with 80 HF samples; Case 2: 64×64 LF (300 samples) with 60 HF samples.
6.2 Reduced High-Fidelity Data Requirement through Time-Derivative Guided Sampling
The time-derivative guided sampling strategy demonstrates remarkable effectiveness, achieving superior predictive accuracy compared to conventional uniform sampling when using the same amount of HF data (Section 6.1). Therefore, in this section, we investigate how this approach can reduce the required HF dataset size while maintaining comparable accuracy to models trained without time-derivative sampling—quantifying the potential data efficiency gains enabled by our physics-guided sampling method. For MF-DeepONet using 64×64 LF resolution data (Case 2), our sampling approach achieves an MSE of 9.008 with only 60 HF samples, surpassing the performance of models trained on the full 100 HF samples with uniform sampling (MSE 9.057, marked as gray in Table 4). This achievement is particularly significant as it represents both improved accuracy and a 40% reduction in required HF data. Similarly, with 32×32 LF resolution data (Case 1), our time-based sampling achieves prediction accuracy (MSE 10.581) using only 80 HF samples, effectively matching the performance of training with the complete HF dataset and uniform sampling (MSE 10.536, marked as gray in Table 4).
Figure 13 provides a visual comparison between the uniform-sampling-based multi-fidelity model trained with the 100 HF dataset and the multi-fidelity model coupled with time-derivative guided sampling approach using only 60 HF dataset. At both time steps ($t=0.5s$ and $t=0.75s$), we observe that both models produce visually comparable flow field predictions that closely match the ground truth. The error distributions demonstrate that, despite using significantly less HF data (only 60%), the time-derivative sampling approach maintains similar prediction quality to the uniform-sampling model. While there are subtle differences in the error patterns—particularly visible in certain regions at $t=0.75s$ where the proposed physics-guided sampling approach shows slightly higher localized errors—these minor differences in error patterns are insignificant when considering the substantial 40% reduction in HF data requirements achieved by our sampling approach. This visual evidence reinforces our quantitative findings that strategic selection of high-fidelity training points based on temporal dynamics enables efficient utilization of limited HF data with minimal degradation in prediction quality.
@@ 6.2 Reduced High-Fidelity Data Requirement through Time-Derivative Guided Sampling
Figure 13: Visual comparison of the predicted vorticity field at $t=0.5s$ and $t=0.75s$ between two models: 1) baseline uniform-sampling-based MF-DeepONet with LF 64×64 and 100 HF samples, and 2) MF-DeepONet with LF 64×64 and only 60 HF samples guided by time-derivative sampling. Despite using 40% fewer HF samples, the model with time-derivative sampling maintains comparable prediction quality, with only minor differences in error patterns.
@@
Figure 14 presents an analysis of error behavior of the test data over time for Cases 1 and 2, comparing scenarios with and without time-derivative guided sampling. It is clear from the plot that incorporating time-derivative guided sampling significantly mitigates error increase as temporal prediction progresses for both cases. This highlights that the physics-guided, time-derivative sampling not only enhances accuracy at individual time steps but also substantially limits error propagation over extended prediction intervals. Such improved long-term predictive stability is particularly valuable for industrial applications where accuracy over prolonged simulation times is crucial.
@@
Figure 14: Comparison of MSE behavior over time steps (200 steps stands for a second) for Case 1 and Case 2, with and without time-derivative guided sampling. The results clearly demonstrate that the time-derivative guided sampling effectively reduces error over time, thereby significantly improving prediction stability and accuracy.
@@
Note: While our sampling strategy leverages temporal derivatives as the key physical indicator, the MF-DeepONet framework itself is not confined to this specific sampling approach. Temporal derivatives serve as an effective physics-informed criterion for identifying critical regions in the examples studied here, capturing fundamental characteristics where rapid changes often indicate important physical transitions. However, the modular nature of our framework allows for integration with various alternative sampling techniques tailored to specific applications. For instance, future work could explore uncertainty-driven sampling for stochastic systems. This flexibility represents a significant advantage of our approach—the underlying MF-DeepONet architecture can be coupled with domain-specific sampling strategies while maintaining its core multi-fidelity benefits. This opens numerous avenues for researchers to develop specialized sampling techniques optimized for their particular physical systems, potentially yielding even greater efficiency gains beyond what we’ve demonstrated with temporal derivatives in this study.

### v2

@@ 5.2 Performance Analysis with Limited High-Fidelity Data
The results shown in Figure 11 reveal several important insights about the MF-DeepONet’s performance with limited HF data. First, examining the MSE trends (Figure 11(a)), we observe a consistent pattern of improved prediction accuracy as the number of HF samples increases across all configurations. The 64×64 resolution configurations consistently outperform their 32×32 counterparts, with the 64×64 with 300 samples combination achieving the best performance across all HF dataset sizes. The training time analysis (Figure 11(b)) reveals an expectable trade-off. Configurations with 300 LF samples consistently require longer training times compared to their 200 LF sample counterparts due to the longer training time required in LF deepONet. However, this increased computational cost appears justified by the improved accuracy, particularly for the 64×64 resolution cases.
@@
Figure 11: MF-DeepONet performance with limited high-fidelity data: different resolutions of LF data with different LF data sizes are explored. The label “LF 32×32 (200)” represents a case where LF data is from a 32×32 resolution with a dataset size of 200—all other labels follow the same notation.
6 Physics-Guided Subsampling for Enhancing Multi-Fidelity DeepONet
6.1 Effects of Physics-Guided Subsampling
Based on our analysis in Section 5.2, we will further investigate two configurations of our fine-tuned MF-DeepONet framework (highlighted in Figure 11(a) as yellow circles with black dashed edges) that achieve an moderate balance between prediction accuracy and high-fidelity data efficiency:
@@ 6.1 Effects of Physics-Guided Subsampling
    Case 1: MF-DeepONet with 32×32 LF resolution (300 samples) and 80 HF samples
@@
    Case 2: MF-DeepONet with 64×64 LF resolution (300 samples) and 60 HF samples
@@
For these two configurations, we analyze the effectiveness of our physics-guided (specifically, time-derivative guided) sampling strategy in practice. This physics-guided approach, formalized in Algorithm 1 in Section 2.2.2, uses a pre-trained LF DeepONet to identify regions with pronounced temporal dynamics. The purpose of this strategy is to make the training of the HF-DeepONet efficient by guiding the selection of a sparse but physically informative subset of spatial points from the full, already available HF simulation dataset, thereby enhancing data utilization efficiency during model training.
In our experiments, we first analyze the effectiveness of our proposed time-derivative guided subsampling. For a more comprehensive comparison, we also present the results from a conventional residual-based sampling strategy, which is a well-established technique in the domain of physics-informed neural networks [27]. This method operates by identifying and sampling points where the errors (residuals) between the LF model’s predictions and the HF ground truth data are large. To ensure a fair comparison between these two guided approaches, we set the sampling ratio to $r=0.1$ for both, meaning that 10% of the HF spatial points are strategically selected based on the respective criterion, while the remaining 90% are sampled uniformly at random to maintain spatial coverage.
The results, detailed in Table 5, demonstrate that while both methods significantly improve accuracy over the baseline, our proposed time-derivative subsampling is slightly superior in both accuracy and computational cost. Our physics-guided approach achieves MSE reductions of 10.02% for Case 1 and 20.73% for Case 2. In comparison, the residual-based method yields slightly smaller MSE reductions of 9.61% and 19.74% for the same cases. Furthermore, our approach is more computationally efficient, with an 8-11% increase in training time overhead, compared to the more significant 18% increase required for residual sampling. This additional overhead for both methods is primarily attributed to the calculations required to guide the sampling (i.e., automatic differentiation for temporal derivatives or residual computation). The greater improvement seen in the Case 2 configurations for both methods indicates that a higher-resolution LF data produces a more accurate LF DeepONet model, which in turn provides more reliable guidance.
@@
Table 5: Performance comparison of different sampling strategies: uniform (Before), proposed time-derivative guided (Time-Deriv. Sampling), and conventional residual-based (Residual).
Metric Case Before Time-Deriv. Sampling Residual Sampling
MSE Case 1: LF 32×32 w/ 80 HF 11.759 10.581 (-10.02%) 10.629 (-9.61%)
Case 2: LF 64×64 w/ 60 HF 11.364 9.008 (-20.73%) 9.121 (-19.74%)
Training hour Case 1: LF 32×32 w/ 80 HF 0.469 0.505 (+7.68%) 0.555 (+18.34%)
Case 2: LF 64×64 w/ 60 HF 0.475 0.526 (+10.74%) 0.562 (+18.32%)
@@
The above quantitative improvements are further supported by the qualitative analysis shown in Figure 12, which compares vorticity profiles at different time steps. At $t=0.25s$ (Figure 12(a)), all models show similar prediction capabilities, closely following the ground truth profile. This comparable performance in the early stage is expected, as the flow field remains relatively simple and closer to the initial conditions. However, at $t=0.5s$ (Figure 12(b)), clear differences emerge in the models’ ability to capture complex flow dynamics. This later time step exhibits notably more complicated flow development with pronounced temporal variations. Indeed, particularly in the middle region ($y=0.3m$ to $0.7m$), models trained with time-derivative subsampling (presented as dashed lines) more accurately reproduce the local fluctuations and subtle variations in the vorticity profile. This superior performance demonstrates that our approach automatically identifies and focuses computational resources on regions with significant temporal dynamics, enhancing prediction accuracy exactly where HF DeepONet should focus. In contrast, models without sampling tend to oversimplify these local features, capturing only the overall trend while missing important flow structure details. These results confirm that our physics-guided subsampling strategy effectively targets time-sensitive regions of the flow field ($t=0.5s$ than $t=0.25s$), showing particular advantage in predicting complex time-dependent dynamics.
@@
Figure 12: Comparison of vorticity profiles from ground truth (GT) and MF-DeepONet predictions with and without time-derivative guided subsampling. At early stages ($t=0.25s$), all models show comparable performance, but at later times ($t=0.5s$), models with time-derivative subsampling demonstrate superior accuracy in capturing complex flow dynamics. Case 1: 32×32 LF (300 samples) with 80 HF samples; Case 2: 64×64 LF (300 samples) with 60 HF samples.
6.2 Reduced High-Fidelity Data Requirement through Physics-Guided Subsampling
The time-derivative guided subsampling strategy demonstrates remarkable effectiveness, achieving superior predictive accuracy compared to conventional uniform sampling when using the same amount of HF data (Section 6.1). Therefore, in this section, we investigate how this approach can reduce the required HF dataset size while maintaining comparable accuracy to models trained without time-derivative subsampling—quantifying the potential data efficiency gains enabled by our physics-guided subsampling method. For MF-DeepONet using 64×64 LF resolution data (Case 2), our sampling approach achieves an MSE of 9.008 with only 60 HF samples, surpassing the performance of models trained on the full 100 HF samples with uniform sampling (MSE 9.057, marked as gray in Table 4). This achievement is particularly significant as it represents both improved accuracy and a 40% reduction in required HF data. Similarly, with 32×32 LF resolution data (Case 1), our time-based sampling achieves prediction accuracy (MSE 10.581) using only 80 HF samples, effectively matching the performance of training with the complete HF dataset and uniform sampling (MSE 10.536, marked as gray in Table 4).
Figure 13 provides a visual comparison between the uniform-sampling-based multi-fidelity model trained with the 100 HF dataset and the multi-fidelity model coupled with time-derivative guided subsampling approach using only 60 HF dataset. At both time steps ($t=0.5s$ and $t=0.75s$), we observe that both models produce visually comparable flow field predictions that closely match the ground truth. The error distributions demonstrate that, despite using significantly less HF data (only 60%), the time-derivative subsampling approach maintains similar prediction quality to the uniform-sampling model. While there are subtle differences in the error patterns—particularly visible in certain regions at $t=0.75s$ where the proposed physics-guided subsampling approach shows slightly higher localized errors—these minor differences in error patterns are insignificant when considering the substantial 40% reduction in HF data requirements achieved by our sampling approach. This visual evidence reinforces our quantitative findings that strategic selection of high-fidelity training points based on temporal dynamics enables efficient utilization of limited HF data with minimal degradation in prediction quality.
@@ 6.2 Reduced High-Fidelity Data Requirement through Physics-Guided Subsampling
Figure 13: Visual comparison of the predicted vorticity field at $t=0.5s$ and $t=0.75s$ between two models: 1) baseline uniform-sampling-based MF-DeepONet with LF 64×64 and 100 HF samples, and 2) MF-DeepONet with LF 64×64 and only 60 HF samples guided by time-derivative subsampling. Despite using 40% fewer HF samples, the model with time-derivative subsampling maintains comparable prediction quality, with only minor differences in error patterns.
@@
Figure 14 presents an analysis of error behavior of the test data over time for Cases 1 and 2, comparing scenarios with and without time-derivative guided subsampling. It is clear from the plot that incorporating time-derivative guided subsampling mitigates error increase as temporal prediction progresses for both cases. This highlights that the physics-guided, time-derivative subsampling not only enhances accuracy at individual time steps but also substantially limits error propagation over extended prediction intervals. Such improved long-term predictive stability is particularly valuable for industrial applications where accuracy over prolonged simulation times is crucial.
@@
Figure 14: Comparison of MSE behavior over time steps (200 steps stands for a second) for Case 1 and Case 2, with and without time-derivative guided subsampling. The results clearly demonstrate that the time-derivative guided subsampling effectively reduces error over time, thereby significantly improving prediction stability and accuracy.
@@
Note on the extensibility of the proposed physics-guided subsampling approach: While our sampling strategy leverages temporal derivatives as the key physical indicator, the MF-DeepONet framework itself is not confined to this specific sampling approach. Temporal derivatives serve as an effective physics-informed criterion for identifying critical regions in the examples studied here, capturing fundamental characteristics where rapid changes often indicate important physical transitions. However, the modular nature of our framework allows for integration with various alternative sampling techniques tailored to specific applications. For instance, future work could explore uncertainty-driven sampling for stochastic systems. This flexibility represents a clear advantage of our approach—the underlying MF-DeepONet architecture can be coupled with domain-specific sampling strategies while maintaining its core multi-fidelity benefits. This opens numerous avenues for researchers to develop specialized sampling techniques optimized for their particular physical systems, potentially yielding even greater efficiency gains beyond what we’ve demonstrated with temporal derivatives in this study.
7 Further Validation with More Challenging Datasets
To further consolidate the robustness and superior performance of our proposed MF-DeepONet framework, particularly in scenarios demanding higher precision and involving more local flow dynamics, we introduce a new set of more challenging datasets. These datasets are designed to incorporate higher spatial resolutions and more complex initial conditions compared to those presented in Section 3. This section first validates our MF-DeepONet framework against conventional approaches using new datasets. Subsequently, we conduct a detailed investigation of our proposed time-derivative method, comparing it with a conventional residual-based approach to assess its efficacy under these more demanding conditions.
7.1 Generation of Higher-Resolution and More Complex Flow Field Datasets
The generation process for these new, more challenging datasets largely follows the methodology outlined in Section 3, utilizing the same pseudo-spectral solver for the incompressible Navier-Stokes equations in a two-dimensional periodic domain $[0,1m]\times[0,1m]$ with a Reynolds number of $Re=1,000$. However, two key aspects are modified:
  1. 1.
    More Complex Initial Conditions: The parameters $var_{1}$ and $var_{2}$ defining the initial velocity field components (Eqn. 6) are now sampled from a uniform distribution in the range $\mathcal{U}[5,7]$ (previously $\mathcal{U}[1,3]$). This change introduces higher spatial frequencies into the initial state of the flow, leading to more complex and rapidly evolving flow structures.
  2. 2.
    Higher Spatial Resolutions: The fidelity levels are increased accordingly as flow characteristics become complicated. For the HF dataset, we now generate 100 training samples and 50 test samples on a finer uniform grid of 256×256 resolution (previously 128×128). For the LF datasets, we generate 300 training samples for each of three upgraded resolutions: 32×32, 64×64, and 128×128 (previously 16×16, 32×32, and 64×64).
These modifications therefore increase the computational cost of data generation. While HF 256×256 data requires approximately 1800 seconds per sample, the LF data requires 75, 120, and 330 seconds for 32×32, 64×64, and 128×128 resolutions, respectively. In total, generating the HF dataset (100 training samples) requires 50 hours, whereas generating each LF dataset (300 samples) requires 6.25, 10, and 27.5 hours, respectively.
The vorticity fields generated using these more challenging conditions are visualized across different resolutions in Figure 15: this figure clearly illustrates the increased complexity compared to the initial dataset presented in Figure 6. The higher values used for $var_{1}$ and $var_{2}$ in the initial conditions (Eqn. 6) result in a greater number of initial vorticity structures and consequently, a more intricate mixing process as the flow evolves. Similar to observations with the previous dataset, Figure 15 shows that the lowest LF resolution (32×32) struggles to represent the detailed flow structures visible in the HF (256×256) simulation, particularly at later time steps. In contrast, the intermediate LF resolutions (64×64 and 128×128) capture the essential macroscopic flow characteristics faithfully.
Figure 15: Effects of grid resolution on vorticity field evolution with fixed parameters $var_{1}=6.4$ and $var_{2}=6.5$ (contours represent vorticity values).
7.2 Proposed MF-DeepONet Framework Performance
Here, we validate the performance of our proposed MF-DeepONet framework against the conventional approach by Lu et al. [29]. Since our initial results in Section 5.1 conclusively showed that our fine-tuning strategy was superior to both full-tuning and linear probing, we focus this validation on our best-performing method (fine-tuning) to ensure a direct comparison against the primary baseline (Lu et al. [29]). The experimental setup, including network architectures, largely mirrors that described in Section 5.1. However, to better leverage the more complex flow physics, the LF DeepONet is now trained for 800 epochs, while the subsequent HF fine-tuning phase remains at 400 epochs.
The results, presented in Table 6, again confirm the robustness of our proposed framework under these more demanding conditions. It is evident that the conventional approach by Lu et al. [29] struggles to achieve a reasonable prediction error across all its configurations, indicating potential limitations when applied to such complex flow physics. In contrast, our proposed MF-DeepONet framework demonstrates much better performance, particularly when leveraging higher-quality LF data. Specifically, the four configurations using LF data at 64×64 and 128×128 resolutions with larger dataset sizes (200 and 300 samples) yield the most accurate and reasonable results. This highlights that the success of our framework is pronounced when the pre-training phase is supported by LF datasets that are both reasonably accurate in resolution and sufficient in quantity, enabling the model to effectively learn and transfer essential flow characteristics for high-fidelity prediction.
Table 6: Comparison of different MF-DeepONet architectures for new dataset. Single-fidelity results are shown for reference (e.g., “Single-fidelity (32×32, 300)” indicates 32×32 resolution with 300 samples).
Architecture LF Data Resolution LF Data Size Training Hour MSE
Single-fidelity (32×32, 300) 1.313 165.292
Single-fidelity (64×64, 300) 1.347 414.936
Single-fidelity (128×128, 300) 1.408 45.075
Single-fidelity (256×256, 100) 0.430 15.918
Lu et al. [29] (conventional MF-DeepONet) 32×32 50 0.498 50.228
100 0.550 60.736
200 0.770 50.059
300 0.976 55.436
64×64 50 0.447 51.301
100 0.564 52.415
200 0.831 50.653
300 1.019 49.644
128×128 50 0.469 50.367
100 0.605 48.326
200 0.809 39.131
300 1.054 32.995
Proposed MF-DeepONet 32×32 50 0.598 29.997
100 0.686 25.242
200 0.915 23.959
300 1.117 20.610
64×64 50 0.576 30.048
100 0.715 23.012
200 0.885 13.919
300 1.058 12.028
128×128 50 0.586 23.745
100 0.708 20.957
200 0.914 18.302
300 1.111 12.364
7.3 Comparative Analysis of Physics-Guided Subsampling Strategy
We further investigate the efficacy of our proposed time-derivative guided subsampling strategy using these new challenging datasets, benchmarking it against the conventional residual-based sampling approach [27]. We apply these two subsampling methods to the four most accurate configurations identified in Table 6 (highlighted in bold). As with our previous sampling experiments, we use a strategic sampling ratio of $r=0.1$.
The performance comparison of these sampling strategies is presented in Table 7. While both subsampling methods offer comparable performance when using the 64x64 LF data, our physics-guided time-derivative strategy demonstrates a clear advantage when a higher-quality LF model is employed. Notably, for the case using 128x128 LF data with 300 samples, time-derivative subsampling yields a 14.20% error reduction, which is more effective than the 11.63% reduction from the residual-based method, while both have similar computational overhead. When the guiding LF model is of moderate quality (e.g., 64x64), its predictions contain more pronounced inaccuracies, making the error-driven residual sampling an effective corrective strategy. However, as the LF model becomes more accurate (e.g., 128x128), its ability to reliably identify physically dynamic regions becomes a more powerful form of guidance. In this regime, focusing on these areas of high temporal activity proves to be a more effective strategy for capturing complex dynamics than simply targeting the largest remaining (and now smaller) errors.
Table 7: Performance comparison of different sampling strategies on the new datasets. "Before" refers to training with uniform subsampling.
Metric Case Before Time-Deriv. Sampling Residual Sampling
MSE LF 64x64 w/ 200 samples 13.919 14.484 (+4.06%) 14.482 (+4.04%)
LF 64x64 w/ 300 samples 12.028 11.929 (-0.82%) 11.889 (-1.16%)
LF 128x128 w/ 200 samples 18.302 18.147 (-0.85%) 18.386 (+0.46%)
LF 128x128 w/ 300 samples 12.364 10.609 (-14.20%) 10.926 (-11.63%)
Training hour LF 64x64 w/ 200 samples 0.885 1.063 (+20.11%) 1.069 (+20.79%)
LF 64x64 w/ 300 samples 1.058 1.279 (+20.89%) 1.276 (+20.60%)
LF 128x128 w/ 200 samples 0.914 1.097 (+19.91%) 1.107 (+21.11%)
LF 128x128 w/ 300 samples 1.111 1.331 (+19.80%) 1.337 (+20.34%)